OpenStudy (anonymous):

Please help Factor 3x2 - 12x - 36 completely.

jimthompson5910 (jim_thompson5910):

3x^2 - 12x - 36 = 3(x^2 - 4x - 12) to factor x^2 - 4x - 12, you need to find two numbers that multiply to -12 and add to -4

OpenStudy (anonymous):

So would it be 3(x+4)(x-3) ?

jimthompson5910 (jim_thompson5910):

4 times -3 is -12 but 4 plus -3 is not -4

OpenStudy (anonymous):

I'm a little confused.

jimthompson5910 (jim_thompson5910):

which factors of -12 also add to -4 ?

OpenStudy (anonymous):

oh 6 and 2?

jimthompson5910 (jim_thompson5910):

example: -1 and 12 multiply to -12 but they don't add to -4

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

Im just confused because my options are 3(x-6)(x+2) , (3x+6)(x-2) , 3(x+4)(x-3) and (3x-4)(x+3) so I dont understand if it 3 and 4 dont work into that and 6 and 2 dont either?

jimthompson5910 (jim_thompson5910):

well it's not 6 and -2 since 6 plus -2 = +4 it's actually -6 and 2 because -6 plus 2 = -4

jimthompson5910 (jim_thompson5910):

so x^2 - 4x - 12 factors to (x-6)(x+2)

jimthompson5910 (jim_thompson5910):

3x^2 - 12x - 36 = 3(x^2 - 4x - 12) 3x^2 - 12x - 36 = 3(x-6)(x+2)

OpenStudy (anonymous):

Oh okay that makes more sense

OpenStudy (anonymous):

could you help me with another question

jimthompson5910 (jim_thompson5910):

sure